can some one help me please
a black racer snake travels 9.2 km in 4 hours. what is the speed in km/h?
I just don't get it. thank you for your time

Found 2 solutions by Earlsdon, richard1234:
Answer by Earlsdon(6294)

If the snake travels 9.2 km in 4 hours, in 1 hour it will travel 4 times less.
So you would divide the 9.2 km by 4 to get 2.3 km/hr.

Answer by richard1234(7193)

Just use (distance) = (rate)*(time)

(rate) = (distance)/(time) = (9.2 km)/(4 hr) = 2.3 km/hr